# Quantum Pufferfish Privacy: A Flexible Privacy Framework for Quantum Systems

### v1

@@ VII-B Formal Guarantees for Auditing QDP
In the previous subsection, we presented possible avenues for auditing QDP, using some tools available in quantum information theory, as well as variational quantum algorithms. Now, we put forward the question “Can we quantify the success of a privacy auditing pipeline for correctly accepting and rejecting an algorithm with given privacy requirements?” The authors of [53, 54] have worked towards answering this question for auditing classical DP, but using a relaxed privacy definition.
In what follows, we propose a hypothesis testing based auditing pipeline for QDP with domain knowledge, which is represented in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$. To obtain a desired significance in the hypothesis test, we employ the trace-norm estimation quantum algorithm proposed in [59], which has at most $\alpha$ additive error from the exact value, with high probability.
@@
Let the estimates of $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})$ and $\mathsf{T}^{\varepsilon}(\sigma,\rho,\mathcal{A})$ from a randomized algorithm (in our analysis we use the algorithm corresponding to [59, Corollary 3.4]) be $\hat{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})$ and $\hat{T}^{\varepsilon}(\sigma,\rho,\mathcal{A})$, respectively. We choose
$\hat{T}(\rho,\sigma,\mathcal{A},\varepsilon)\coloneqq\max\{\hat{T}^{\varepsilon}(\rho,\sigma,\mathcal{A}),\hat{T}^{\varepsilon}(\sigma,\rho,\mathcal{A})\}$ (212)
@@
Lemma 3 (Estimating trace distance using samples of $\rho$, $\sigma$ — restatement of [59, Corollary 3.4]).
Given access to identical copies of $d$-dimensional quantum states $\rho$ and $\sigma$, there is a quantum algorithm that estimates $\mathsf{T}(\rho,\sigma)$ within additive error $\alpha$ and with probability not less than $(1-\beta)$, by using
@@
where $\mathrm{sgn}(V)\coloneqq W\mathrm{sgn}(\Sigma)U^{\dagger}$, and $\mathrm{sgn}(\cdot)$ is the sign function. Then, $\mathrm{Tr}\!\left[\rho\,\mathrm{sgn}(V)\right]$ and $\mathrm{Tr}\!\left[\sigma\,\mathrm{sgn}(V)\right]$ can be estimated separately, by combining the techniques of quantum singular value transformation [62] and the Hadamard test [63]. To this end, to implement unitary block-encodings of $\rho$ and $\sigma$ approximately, the technique of density matrix exponentiation [64] is used. The same techniques can be employed to compute $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})$ since
@@
$\sup_{\rho\sim\sigma}\mathbb{P}\!\left(\hat{T}(\rho,\sigma,\mathcal{A},\varepsilon)>\mathsf{g}(\varepsilon,\delta)+\alpha\,\middle|\,H_{0}\right)\leq\beta,$ (216)
@@
Fix $\rho$ and $\sigma$ such that $\rho\sim\sigma$. Under the null hypothesis and the assumption that $\hat{T}(\rho,\sigma,\mathcal{A},\varepsilon)={\hat{T}}^{\varepsilon}(\rho,\sigma,\mathcal{A})$, we have that
$\mathbb{P}\!\left(\hat{T}(\rho,\sigma,\mathcal{A},\varepsilon)>\mathsf{g}(\varepsilon,\delta)+\alpha\right)$
$=\mathbb{P}\!\left({\hat{T}}^{\varepsilon}(\rho,\sigma,\mathcal{A})>\mathsf{g}(\varepsilon,\delta)+\alpha\right)$ (219)
$\stackrel{{\scriptstyle(a)}}{{\leq}}\mathbb{P}\!\left({\hat{T}}^{\varepsilon}(\rho,\sigma,\mathcal{A})-{{\mathsf{T}}}^{\varepsilon}(\rho,\sigma,\mathcal{A})>\alpha\right)$ (220)
$\leq\mathbb{P}\!\left(|{\hat{T}}^{\varepsilon}(\rho,\sigma,\mathcal{A})-{{\mathsf{T}}}^{\varepsilon}(\rho,\sigma,\mathcal{A})|>\alpha\right)$ (221)
@@
where: (a) follows since ${{\mathsf{T}}}^{\varepsilon}(\rho,\sigma,\mathcal{A})\leq\mathsf{g}(\varepsilon,\delta)$ under the null hypothesis; (b) from the high probability statement in Lemma 3. Similarly, the above inequality holds when $\hat{T}(\rho,\sigma,\mathcal{A},\varepsilon)={\hat{T}}^{\varepsilon}(\sigma,\rho,\mathcal{A})$ concluding the proof. ∎
@@
Remark 15 (Computational complexity with the rank $r$).
From Proposition 9, it is evident that the copy complexity of the algorithm grows as $O(r^{2}\log^{2}(r))$, where $r$ is the rank upper bound. Let $\mathcal{A}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ be a quantum channel. Then, $r\leq d_{B}$, where $d_{B}$ is the dimension of the Hilbert space $\mathcal{H}_{B}$. Considering computational complexity, one possibility is to compose $\mathcal{A}$ with another quantum channel $\mathcal{N}$ that translates the space to a low-dimensional setting. However, due to the data-processing inequality for the trace distance, it will only provide a lower bound. With that, it is possible to reject algorithms if $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{N}\circ\mathcal{A})>g(\varepsilon,\delta)$, which implies that $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})>g(\varepsilon,\delta)$. This approach will then have the limitations equipped with the classical auditing approaches using relaxed privacy notions, since the contraction gap between $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{N}\circ\mathcal{A})$, and $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})$ is hard to quantify. In the quantification of the gap, finding the contraction coefficient $\eta_{\mathcal{N}}$ of the channel $\mathcal{N}$ would be useful if $\eta_{\mathcal{N}}<1$.
In summary, we presented in this section a hypothesis testing pipeline for auditing the privacy of quantum systems while providing formal guarantees on auditing QDP using quantum algorithms designed for trace distance. Analysing the Type-II error of this approach is also an important task to quantify the power of the test, which is left for future work.
@@ VIII Information-Theoretic Bounds from QPP and Connections to Quantum Fairness
In this section, we first explore several information-theoretic bounds that arise from an algorithm satisfying QPP constraints. We then demonstrate how these bounds can be used to establish connections between QPP and quantum fairness [25, 24]. Later on we utilize the results derived in this section to determine the relative strength of the QPP variants introduced in Section IX.
@@ VIII-A Information-Theoretic Bounds from QPP
It was pointed out in [10] that finding bounds on quantum relative entropy and mutual information due to QDP is an interesting open problem. In this section, we address this question in a general setting, which includes QDP as a special case, by giving bounds for relative entropy and Holevo information. We additionally provide bounds in terms of Rényi relative entropies, and trace distance. For the rest of the discussion, fix the privacy framework to be $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$.
@@
Fix $\alpha>1$. If $\mathcal{A}$ is $\varepsilon$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$ for all $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$, then
@@
where $\mathsf{D}_{T}$ is the Thompson metric from (17). By this implication and the fact that every quantum Rényi-divergence $\mathsf{D}_{\alpha}$ of order $\alpha$ satisfying data processing is bounded from above by $\mathsf{D}_{\max}$ [65, Eq. (4.36)], $\varepsilon$-QPP implies that for all $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$,
@@
By a different argument via the maximal extension presented in Lemma 4, we can further arrive at the following, for all $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$,
@@
Fix $\alpha>1$, and $\rho,\sigma$ PSD operators. The following inequality holds:
$\mathsf{D}_{\alpha}(\rho\|\sigma)\leq\frac{\alpha}{2}(\mathsf{D}_{T}(\rho\|\sigma))^{2}.$ (228)
@@
An operational interpretation of the Thompson metric has appeared in symmetric postselected hypothesis testing (a setting allowing for an inconclusive outcome along with two general conclusive outcomes and postselecting on the conclusive outcomes) as the asymptotic error exponent of discriminating two quantum states $\rho$ and $\sigma$ [41], as well as in the resource theory of symmetric distinguishability [40]. Here by referring to (225), the QPP framework also provides another operational interpretation of the Thompson metric. In the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$, for fixed $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$ (by definition also for $(\mathcal{T},\mathcal{R})\in\mathcal{Q}$), the Thompson metric given by $\mathsf{D}_{T}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right)$ is equal to the minimal $\varepsilon$ needed to achieve $\varepsilon$-QPP for the secret pair $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$.
@@
$\sup_{\Theta,(\mathcal{R},\mathcal{T})\in\mathcal{Q}}\left\|\mathcal{A}(\rho^{\mathcal{R}})-\mathcal{A}(\rho^{\mathcal{T}})\right\|_{1}\leq 1-\frac{2(1-\delta)}{e^{\varepsilon}+1}.$ (230)
@@
The first inequality holds by applying the quantum Pinsker inequality ($\frac{1}{2}\left\|\rho-\sigma\right\|^{2}_{1}\leq\mathsf{D}(\rho\|\sigma)$) [66, Theorem 1.15] and Proposition 10.
For the second inequality, next we prove a more general relation: $(\varepsilon,\delta)$-QPP implies $(\varepsilon^{\prime},\delta^{\prime})$-QPP for $\varepsilon^{\prime}<\varepsilon$ with
$\delta^{\prime}=1-(e^{\varepsilon^{\prime}}+1)(1-\delta)/(e^{\varepsilon}+1),$ (231)
similar to [6, Property 3] for classical DP. Using that, fixing $\varepsilon^{\prime}$ therein to zero leads to the desired result.
Proof that $(\varepsilon,\delta)$-QPP implies $(\varepsilon^{\prime},\delta^{\prime})$-QPP:
With the above choice for $\delta^{\prime}$, we have $\delta-\delta^{\prime}=(1-\delta)(e^{\varepsilon^{\prime}}-e^{\varepsilon})$. Plugging this in, we find that
$\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{R}})\right]$
By choosing the measurement operator $\mathrm{I}-\mathrm{M}$, we also have
Rewriting (235), we arrive at
Similar to the previous manipulations, we get
$\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{R}})\right]\leq e^{\varepsilon^{\prime}}\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]+\delta^{\prime}\\ +(e^{\varepsilon^{\prime}}-e^{-\varepsilon})\left(-\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]+\frac{1-\delta}{e^{\varepsilon}+1}\right).$ (237)
A similar inequality holds for every $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$ and $P_{X}\in\Theta$. Thus, the desired implication has been proved. ∎
Bounds on mutual information
Proposition 12 (Bounds on Holevo information due to QLDP).
Furthermore, if $\mathcal{A}$ is $(\varepsilon,\delta)$-QLDP, then
is the binary Shannon entropy in nats.
@@
The following proof ideas are inspired by the works on mutual information based classical DP and PP presented in [6, Theorem 1] and [54, Theorem 1], respectively. For the proof of the first part, consider that
$\mathsf{I}\!\left(X;B\right)_{\sigma}$
$\leq\min\!\left\{\varepsilon,\frac{\varepsilon^{2}}{2}\right\},$ (245)
For the second part, consider that
$\leq\delta^{\prime}\ln(d-1)+\mathsf{h}(\delta^{\prime}),$ (247)
where
$\bar{\rho}=\sum_{x\in\mathcal{X}}P_{X}(x)\rho^{x},$ (248)
and the final inequality follows from the fact that $\left\|\mathcal{A}(\bar{\rho})-\mathcal{A}(\rho^{x})\right\|_{1}\leq\delta^{\prime}$ due to $\mathcal{A}$ satisfying $(\varepsilon,\delta)$-QPP along with Proposition 11. Then, to arrive at the final expression, we use the entropy continuity result known as the Fannes–Audenaert Inequality [68]:
concluding the proof. ∎
Proposition 13 (Bounds on mutual information due to QDP).
Proof:
$\omega^{x}$ $\coloneqq\rho_{i}^{x}\otimes\rho^{n\backslash i}.$ (255)
For the first part, it thus follows that
For the second part, consider that
$\mathsf{I}(X_{i};B|X^{n\backslash i}=z^{n\backslash i})_{\sigma}$
$\leq\delta^{\prime}\ln(d-1)+\mathsf{h}(\delta^{\prime}),$ (258)
where the last inequality holds because $\left\|\mathcal{A}(\omega)-\mathcal{A}(\omega^{x})\right\|_{1}\leq\delta^{\prime}$ with $\mathcal{A}$ being $(\varepsilon,\delta)$-QDP for the pair $\sigma\sim\sigma^{x}$, and again applying the continuity result for quantum entropy, as in the proof of Proposition 12. ∎
By the data processing inequality for mutual information, we have $\mathsf{I}(X_{i};Y|X^{n\backslash i})\leq\varepsilon^{\prime}$, where $\varepsilon^{\prime}$ is determined by the values on the right-hand side of the inequalities in Proposition 13. This showcases that the setup of Proposition 13 satisfies $\varepsilon^{\prime}$-mutual information differential privacy, as proposed in [6], where a randomized mechanism $A:\mathcal{X}^{n\times k}\to\mathcal{Y}$ is defined to be $\varepsilon$-mutual information differentially private if

### v2

@@ VII-B Formal Guarantees for Auditing QDP
The question of quantifying the success of a privacy auditing approach, specifically in correctly accepting and rejecting an algorithm with given privacy requirements, is a crucial consideration in privacy auditing research. The authors of [68, 10] have worked towards answering this question for auditing classical DP, but using a relaxed privacy definition.
To tackle this for the quantum setting, we propose a hypothesis testing-based auditing pipeline tailored specifically for QDP (also for QPP). In this pipeline, we use the trace-norm estimation quantum algorithm proposed in [73]. This quantum algorithm provides an estimation with at most $\alpha$ additive error from the exact value, with high probability, which allows us to achieve the desired significance in the hypothesis test.
@@
Let the estimates of $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})$ and $\mathsf{T}^{\varepsilon}(\sigma,\rho,\mathcal{A})$ from a randomized algorithm (in our analysis we use the algorithm corresponding to [73, Corollary 3.4]) be $\hat{\mathsf{T}}^{\varepsilon}(\rho,\sigma,\mathcal{A})$ and $\hat{\mathsf{T}}^{\varepsilon}(\sigma,\rho,\mathcal{A})$, respectively. We choose
$\hat{\mathsf{T}}^{\varepsilon}_{\max}(\rho,\sigma,\mathcal{A})\coloneqq\max\{\hat{\mathsf{T}}^{\varepsilon}(\rho,\sigma,\mathcal{A}),\hat{\mathsf{T}}^{\varepsilon}(\sigma,\rho,\mathcal{A})\}$ (185)
@@
Lemma 3 (Estimating trace distance using samples of $\rho$, $\sigma$ — restatement of [73, Corollary 3.4]).
Given access to identical copies of $d$-dimensional quantum states $\rho$ and $\sigma$, there is a quantum algorithm that estimates the normalized trace distance $\mathsf{T}(\rho,\sigma)$ (recall (2)) within additive error $\alpha$ and with probability not less than $1-\beta$, by using
@@
where $\mathrm{sgn}(V)\coloneqq W\mathrm{sgn}(\Sigma)U^{\dagger}$, and $\mathrm{sgn}(\cdot)$ is the sign function. Then, $\mathrm{Tr}\!\left[\rho\,\mathrm{sgn}(V)\right]$ and $\mathrm{Tr}\!\left[\sigma\,\mathrm{sgn}(V)\right]$ can be estimated separately, by combining the techniques of quantum singular value transformation [76] and the Hadamard test [77]. To this end, to implement unitary block-encodings of $\rho$ and $\sigma$ approximately, the technique of density matrix exponentiation [78] is used.
The same techniques can be employed to compute $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})$ since
@@
$\sup_{\rho\sim\sigma}\mathbb{P}\!\left(\hat{\mathsf{T}}^{\varepsilon}_{\max}(\rho,\sigma,\mathcal{A})>\mathsf{g}(\varepsilon,\delta)+\alpha\,\middle|\,H_{0}\right)\leq\beta,$ (189)
@@
Fix $\rho$ and $\sigma$ such that $\rho\sim\sigma$. Under the null hypothesis and the assumption that $\hat{\mathsf{T}}^{\varepsilon}_{\max}(\rho,\sigma,\mathcal{A})={\hat{\mathsf{T}}}^{\varepsilon}(\rho,\sigma,\mathcal{A})$, we have that
$\mathbb{P}\!\left(\hat{\mathsf{T}}^{\varepsilon}_{\max}(\rho,\sigma,\mathcal{A})>\mathsf{g}(\varepsilon,\delta)+\alpha\right)$
$=\mathbb{P}\!\left({\hat{\mathsf{T}}}^{\varepsilon}(\rho,\sigma,\mathcal{A})>\mathsf{g}(\varepsilon,\delta)+\alpha\right)$ (192)
$\stackrel{{\scriptstyle(a)}}{{\leq}}\mathbb{P}\!\left({\hat{\mathsf{T}}}^{\varepsilon}(\rho,\sigma,\mathcal{A})-{{\mathsf{T}}}^{\varepsilon}(\rho,\sigma,\mathcal{A})>\alpha\right)$ (193)
$\leq\mathbb{P}\!\left(|{\hat{\mathsf{T}}}^{\varepsilon}(\rho,\sigma,\mathcal{A})-{{\mathsf{T}}}^{\varepsilon}(\rho,\sigma,\mathcal{A})|>\alpha\right)$ (194)
@@
where: (a) follows since ${{\mathsf{T}}}^{\varepsilon}(\rho,\sigma,\mathcal{A})\leq\mathsf{g}(\varepsilon,\delta)$ under the null hypothesis; (b) from the high probability statement in Lemma 3. Similarly, the above inequality holds when $\hat{\mathsf{T}}^{\varepsilon}_{\max}(\rho,\sigma,\mathcal{A})={\hat{\mathsf{T}}}^{\varepsilon}(\sigma,\rho,\mathcal{A})$ concluding the proof. ∎
@@
Remark 17 (Computational complexity with rank $r$).
From Proposition 8, it is evident that the copy complexity of the algorithm grows as $O(r^{2}\log^{2}(r))$. Let $\mathcal{A}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ be a quantum channel. Then, $r\leq d_{B}$, where $d_{B}$ is the dimension of the Hilbert space $\mathcal{H}_{B}$. To handle computational complexity, one possibility is to compose $\mathcal{A}$ with another quantum channel $\mathcal{N}$ that translates the space to a low-dimensional setting. However, due to the data-processing inequality for the trace distance, it will only provide a lower bound. With that, it is possible to reject algorithms if $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{N}\circ\mathcal{A})>g(\varepsilon,\delta)$, which implies that $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})>g(\varepsilon,\delta)$. Consequently, it may lead to limitations similar to classical auditing approaches that use relaxed privacy notions, since the contraction gap between $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{N}\circ\mathcal{A})$, and $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})$ is hard to quantify. In the quantification of the gap, finding the contraction coefficient $\eta_{\mathcal{N}}$ of the channel $\mathcal{N}$ would be useful if $\eta_{\mathcal{N}}<1$ (recall that the contraction coefficient of a channel $\mathcal{N}$ under a generalized divergence $\boldsymbol{\mathsf{D}}$, as given in Eq. 1, is defined as $\eta_{\mathcal{N}}\coloneqq\sup_{\rho,\sigma\in\mathcal{D},\boldsymbol{\mathsf{D}}(\rho\|\sigma)\neq 0}\frac{\boldsymbol{\mathsf{D}}\left(\mathcal{N}(\rho)\|\mathcal{N}(\sigma)\right)}{\boldsymbol{\mathsf{D}}(\rho\|\sigma)}$).
In summary, we proposed a hypothesis testing pipeline for auditing the privacy of quantum systems, offering formal guarantees on auditing QDP using quantum algorithms designed for estimating trace distance. However, an essential task for further investigation is analyzing the Type-II error of this approach. This analysis would allow us to quantify the power of the test and assess its ability to correctly accept algorithms with the desired privacy requirements, which is left for future work.
@@ VIII Information-Theoretic Bounds from QPP and Connections to Quantum Fairness
In this section, we begin by investigating several information-theoretic bounds that stem from an algorithm satisfying QPP constraints. We then establish connections between QPP and quantum fairness [35, 34] using these bounds. Later on we utilize the derived bounds to assess the relative strength of the QPP variants introduced in Section IX.
@@ VIII-A Information-Theoretic Bounds from QPP
In [15], it was highlighted that finding bounds on quantum relative entropy and mutual information resulting from QDP is an interesting open problem. We address this question in a general setting, encompassing QDP as a special case. We offer bounds for relative entropy and Holevo information, along with bounds for Rényi relative entropies and trace distance. For the rest of the discussion, we adopt the fixed privacy framework to be $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$.
@@
Fix $\alpha>1$. If $\mathcal{A}$ is $\varepsilon$-QPP (i.e., $(\varepsilon,0)$-QPP) in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$ for all $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$, then
@@
where $\mathsf{D}_{T}$ is the Thompson metric from (16). This follows by the definition of the max-relative entropy defined in (14) and Thompson metric in (16), and recalling the definition of QPP framework for $\delta=0$ and $\mathcal{M}=\bar{\mathcal{M}}$ (see Definition 4). By this implication and the fact that every quantum Rényi-divergence $\mathsf{D}_{\alpha}$ of order $\alpha$ satisfying data processing is bounded from above by $\mathsf{D}_{\max}$ [79, Eq. (4.36)], $\varepsilon$-QPP implies that for all $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$,
@@
By a different argument via the maximal extension presented in Lemma 4, we can further arrive at the following, for all $P_{X}\in\Theta$, $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$ and $\varepsilon\alpha<2$,
@@
Fix $\alpha>1$, and $\rho,\sigma$ PSD operators. For $\alpha\mathsf{D}_{T}(\rho\|\sigma)\leq 2$, the following inequality holds:
$\mathsf{D}_{\alpha}(\rho\|\sigma)\leq\frac{\alpha}{2}(\mathsf{D}_{T}(\rho\|\sigma))^{2},$ (201)
where $\mathsf{D}_{\alpha}$ is an arbitrary quantum Rényi relative entropy satisfying data processing.
@@
An operational interpretation of the Thompson metric has appeared in symmetric postselected hypothesis testing (a setting allowing for an inconclusive outcome along with two general conclusive outcomes and postselecting on the conclusive outcomes) as the asymptotic error exponent of discriminating two quantum states $\rho$ and $\sigma$ [51], as well as in the resource theory of symmetric distinguishability [50]. Here by referring to (198), the QPP framework also provides another operational interpretation of the Thompson metric. In the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$, for fixed $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})~{}\in~{}\mathcal{Q}$,999by definition also for $(\mathcal{T},\mathcal{R})\in\mathcal{Q}$ the Thompson metric given by $\mathsf{D}_{T}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right)$ is equal to the minimal $\varepsilon$ needed to achieve $\varepsilon$-QPP.
@@
$\sup_{\Theta,(\mathcal{R},\mathcal{T})\in\mathcal{Q}}\left\|\mathcal{A}(\rho^{\mathcal{R}})-\mathcal{A}(\rho^{\mathcal{T}})\right\|_{1}\leq 2-\frac{4(1-\delta)}{e^{\varepsilon}+1}.$ (203)
@@
The first inequality holds by applying the quantum Pinsker inequality ($\frac{1}{2}\left\|\rho-\sigma\right\|^{2}_{1}\leq\mathsf{D}(\rho\|\sigma)$) [80, Theorem 1.15] and Proposition 9.
For the second inequality: $(0,\delta^{\prime})$-QPP is equivalent to
$\sup_{\Theta,(\mathcal{R},\mathcal{T})\in\mathcal{Q}}\frac{\left\|\mathcal{A}(\rho^{\mathcal{R}})-\mathcal{A}(\rho^{\mathcal{T}})\right\|_{1}}{2}\leq\delta^{\prime}.$ (204)
Then, adapting Lemma 5 and fixing $\varepsilon^{\prime}$ therein to zero leads to the desired result. ∎
Lemma 5.
Fix $(\mathcal{S},\mathcal{Q},\Theta,{\mathcal{M}})$ privacy framework. Then, we have
[…]
where $\varepsilon^{\prime}<\varepsilon$ with
$\delta^{\prime}\coloneqq 1-\frac{(e^{\varepsilon^{\prime}}+1)(1-\delta)}{(e^{\varepsilon}+1)}.$ (206)
@@
The proof follows similarly to the proof of [11, Property 3] for classical DP and is presented in Appendix I. ∎
Remark 19 (Bounds on Holevo information).
In Appendix J, we provide bounds on the Holevo information in the settings of QDP and QLDP (recall this privacy notion from Section III-C2).